A progressive wave is represented by [y=Asin2π(0.5t-0.1x).find
Period
Frequency
Wavelength
Velocity of the wave
1
Expert's answer
2019-04-25T09:22:41-0400

The wave equation is as follows

"y(x,t)=A\sin\left[2\pi\left(\frac{t}{T}-\frac{x}{\lambda}\right)\right]"

In our case

"y(x,t)=A\sin [2\pi(0.5t-0.1x)]"

So, the period


"T=2\:\rm{s}"

The frequency

"f=\frac{1}{T}=\frac{1}{2}=0.5\:\rm{Hz}"

The wavelength

"\lambda=10\:\rm{m}"

The velocity of the wave

"v=\frac{\lambda}{T}=\frac{10\:\rm{m}}{2\:\rm{s}}=5\:\rm{m/s}"